OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length of feet 180 and a width of 125 feet. Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. What type of object in your home or school might this be?

OpenStudy (anonymous):

if all your trying to find is perimeter all you do is add the sides up. 125+125=250 and 180+180=360. 250+360=610, unless I read the problem wrong.